# Customized Interior-Point Methods Solver for Embedded Real-Time Convex Optimization

### v1

@@ VI Numerical Experiments on Embedded System
The solver developed in this study is specifically designed for embedded real-time optimization. Therefore, we evaluated its performance in the embedded system environment. The evaluations were conducted on an AMD Zynq-7000 system-on-chip featuring an ARM Cortex-A9 processor running at a maximum clock frequency of 650MHz with 512MB of RAM. For comparison, we used ECOS and SCS, as both are implemented in C and distributed as open-source software, facilitating deployment on embedded systems. All solvers were compiled with the -O3 optimization level to maximize execution efficiency.
@@ VI-A Trajectory Optimization for Mars Landing
Figure 5: Computation time results for the trajectory optimization examples
@@
$\begin{split}\underset{x,z,u,\sigma}{\text{minimize}}&\quad-z_{N}\\ \text{subject to}&\quad x_{k+1}=Ax_{k}+Bu_{k}+g\quad\forall k\in\{0,1,\cdots,N-1\}\\ &\quad z_{k+1}=z_{k}-\alpha\sigma_{k}\Delta t\quad\forall k\in\{0,1,\cdots,N-1\}\\ &\quad x_{0}=x_{\text{init}},\ z_{0}=\ln(m_{\text{wet}}),\ x_{N}={\bf{0}}_{6\times 1}\\ &\quad\mu_{1,k}\left[1-(z_{k}-z_{0,k})\right]\leq\sigma_{k}\quad\forall k\in\{0,1,\cdots,N\}\\ &\quad\sigma_{k}\leq\mu_{2,k}\left[1-(z_{k}-z_{0,k})\right]\quad\forall k\in\{0,1,\cdots,N\}\\ &\quad\|u_{k}\|\leq\sigma_{k}\quad\forall k\in\{0,1,\cdots,N\}\\ &\quad e_{3}^{T}u_{k}\geq\sigma_{k}\cos{\theta_{\max}}\quad\forall k\in\{0,1,\cdots,N\}\\ &\quad z_{N}\geq\ln(m_{\text{dry}})\end{split}$
@@
where $x_{k}\in\mathbb{R}^{6}$ is the position and velocity of the lander, $z_{k}\in\mathbb{R}$ is the natural logarithm of the lander’s mass, $u_{k}\in\mathbb{R}^{3}$ is the acceleration generated by the lander’s thrusters, and $\sigma_{k}\in\mathbb{R}$ is the slack variable at time node $k$. The parameters $m_{\text{dry}}$ and $m_{\text{wet}}$ represent the dry mass and wet mass of the lander, respectively, $e_{3}\in\mathbb{R}^{3}$ is the $z-$axis basis vector of the landing inertial coordinate frame, $\theta_{\max}$ is the maximum tilt angle of the lander from $e_{3}$ vector, and $x_{\text{init}}\in\mathbb{R}^{6}$ is the initial states of the lander. We define $z_{0,k}=\ln(m_{\text{wet}}-\alpha\rho_{2}k\Delta t)$, $\mu_{1,k}=\rho_{1}e^{-z_{0,k}}$, $\mu_{2,k}=\rho_{2}e^{-z_{0,k}}$ where $\rho_{1}$ and $\rho_{2}$ is the minimum and maximum thrust magnitudes, respectively, $\alpha$ is the fuel depletion parameter, and $\Delta t$ is the time step between nodes. The matrices $A$ and $B$, and the vector $g$ in the dynamics are given by:
$\begin{split}A=\begin{bmatrix}I_{3}&I_{3}\Delta t\\ {\bf{0}}_{3\times 3}&I_{3}\end{bmatrix},\quad B=\begin{bmatrix}(\frac{1}{2}\Delta t^{2})I_{3}\\ I_{3}\Delta t\end{bmatrix}\\ g=\begin{bmatrix}0&0&-\frac{1}{2}g_{0}\Delta t^{2}&0&0&-g_{0}\Delta t\end{bmatrix}^{T}\end{split}$
Ten problem instances were generated by increasing the number of nodes $N$ from 25 to 500, and the following parameter values were used in this numerical experiment.
@@
$\centering\begin{split}x_{\text{init}}=[200.0,\ 0.0,\ 800.0,\ -35.0,\ 0.0,\ -75.0]^{T}\\ \rho_{1}=7440,\ \rho_{2}=18600,\ g_{0}=3.7114,\ \theta_{\max}=\pi/12\\ m_{\text{dry}}=1505,\ m_{\text{wet}}=1905,\ \alpha=4.53\times 10^{-4},\ t_{f}=48.0\end{split}\@add@centering$
We set the termination tolerance to $10^{-8}$ for the developed customized solver and ECOS, while a relaxed tolerance of $10^{-4}$ was used for SCS, considering the convergence characteristics of ADMM.
@@
Results—Figure 5 presents the computation time results for the Mars landing trajectory optimization problems. We can observe that the developed solver and ECOS achieve faster onboard computation times than SCS even under tighter termination tolerances, demonstrating the effectiveness of IPMs for G&C applications. For small-scale instances ($N\leq 150$), ECOS and the developed solver exhibit comparable performance. However, for larger-scale instances, the solver developed in this study demonstrates faster performance. This improvement is attributed to the elimination of unnecessary logic and the enhanced efficiency of data memory management achieved through solver customization.
@@ VI-B Quadcopter Model Predictive Control
Figure 6: Computation time results for the quadcopter MPC examples
@@
$\begin{split}\underset{x,u}{\text{minimize}}&\quad(x_{N}-x_{r})^{T}Q_{N}(x_{N}-x_{r})\ \\ &\quad+\sum_{k=0}^{N-1}\left[(x_{k}-x_{r})^{T}Q(x_{k}-x_{r})+u_{k}^{T}Ru_{k}\right]\\ \text{subject to}&\quad x_{k+1}=Ax_{k}+Bu_{k}\quad\forall k\in\{0,1,\cdots,N-1\}\\ &\quad x_{0}=x_{\text{init}}\\ &\quad x_{\min}\leq x_{k}\leq x_{\max}\quad\forall k\in\{0,1,\cdots,N\}\\ &\quad u_{\min}\leq u_{k}\leq u_{\max}\quad\forall k\in\{0,1,\cdots,N-1\}\end{split}$
@@
$\begin{split}Q=Q_{N}&=\text{diag}(10,10,10,5,5,5,0,0,10,{\bf{0}}_{1\times 3})\\ R&=0.1\times I_{4}\\ x_{\text{init}}&=\begin{bmatrix}1.0&0.0&0.5&{\bf{0}}_{1\times 9}\end{bmatrix}^{T}\\ \quad x_{r}&=\begin{bmatrix}{\bf{0}}_{1\times 6}&0.0&0.0&1.0&{\bf{0}}_{1\times 3}\end{bmatrix}^{T}\\ \end{split}$
@@
Results—Figure 6 presents the computation time results for the quadcopter MPC problems. The results demonstrate that the solver developed in this study outperforms the other solvers. Such performance is mainly attributed to the efficient algorithmic formulation that directly handles quadratic cost functions and the fast convergence properties of IPMs. This capability is particularly important in MPC applications, where the optimization problem should be solved repeatedly in real-time based on feedback from the updated states of the quadcopter.
@@ VII Conclusion
In this work, we have developed a customized IPM-based solver specifically tailored for embedded real-time convex optimization. The developed solver is capable of handling problems within the class of SOCPs, while supporting infeasibility detection through a homogeneous embedding formulation. Unlike conventional homogeneous self-dual embedding approaches, which are limited to problems with linear cost functions, we adopt a homogeneous embedding framework for MCPs from previous work which can directly accommodate quadratic objective terms. This formulation enables efficient and robust algorithm design by avoiding the reformulation of quadratic objectives into conic constraints. To support embedded system deployment, the solver is implemented in C with complete static memory allocation, making it suitable for resource-constrained systems. A code generation tool is also developed to automate the solver customization process for a given problem family, exploiting structural sparsity offline and minimizing computation overhead during online execution. Numerical benchmark and embedded hardware platform experiments demonstrate that the developed solver outperforms existing IPM- and ADMM-based solvers in terms of computation time and reliability. The results highlight its effectiveness for onboard optimization in G&C applications such as MPC and trajectory optimization.
@@ Acknowledgments
This work has been supported by the Korea Aerospace Research Institute (KARI)’s own research project titled “Development of Control Performance Analysis Program for Spaceplanes”. The authors acknowledge the use of ChatGPT to improve the syntax and grammar of several paragraphs in the manuscript. The manuscript was originally drafted by the authors, with the AI system employed to enhance clarity and readability without modifying the technical content.

### v2

@@ VI Numerical Experiments on Embedded System
The solver developed in this study is specifically designed for embedded real-time optimization. Therefore, we evaluated its performance on G&C problems in the embedded system environment. The evaluations were conducted on an AMD Zynq-7000 system-on-chip featuring an ARM Cortex-A9 processor running at a maximum clock frequency of 650MHz with 512MB of RAM. For comparison, we used ECOS and SCS, as both are implemented in C and distributed as open-source software, facilitating deployment on embedded systems. All solvers were compiled with the -O3 optimization level to maximize execution efficiency.
@@ VI-A Trajectory Optimization for Mars Landing
$\begin{split}\underset{x,z,u,\sigma}{\text{minimize}}&\quad-z_{N}\\ \text{subject to}&\quad x_{k+1}=A_{k}x_{k}+B_{k}u_{k}+g\ \forall k\in\{0,1,\cdots,N-1\}\\ &\quad z_{k+1}=z_{k}-\alpha\sigma_{k}\Delta t\quad\forall k\in\{0,1,\cdots,N-1\}\\ &\quad x_{0}=x_{\text{init}},\ z_{0}=\ln(m_{\text{wet}}),\ x_{N}={\bf{0}}_{6\times 1}\\ &\quad\mu_{1,k}\left[1-(z_{k}-z_{0,k})\right]\leq\sigma_{k}\quad\forall k\in\{0,1,\cdots,N\}\\ &\quad\sigma_{k}\leq\mu_{2,k}\left[1-(z_{k}-z_{0,k})\right]\quad\forall k\in\{0,1,\cdots,N\}\\ &\quad\|u_{k}\|\leq\sigma_{k}\quad\forall k\in\{0,1,\cdots,N\}\\ &\quad e_{3}^{T}u_{k}\geq\sigma_{k}\cos{\theta_{\max}}\quad\forall k\in\{0,1,\cdots,N\}\\ &\quad z_{N}\geq\ln(m_{\text{dry}})\end{split}$ (72)
@@
where $x_{k}\in\mathbb{R}^{6}$ is the position and velocity of the lander, $z_{k}\in\mathbb{R}$ is the natural logarithm of the lander’s mass, $u_{k}\in\mathbb{R}^{3}$ is the acceleration generated by the lander’s thrusters, and $\sigma_{k}\in\mathbb{R}$ is the slack variable at time node $k$. The parameters $m_{\text{dry}}$ and $m_{\text{wet}}$ represent the dry mass and wet mass of the lander, respectively, $e_{3}\in\mathbb{R}^{3}$ is the $z$-axis basis vector of the landing inertial coordinate frame, $\theta_{\max}$ is the maximum tilt angle of the lander from $e_{3}$ vector, and $x_{\text{init}}\in\mathbb{R}^{6}$ is the initial states of the lander. We define $z_{0,k}=\ln(m_{\text{wet}}-\alpha\rho_{2}k\Delta t)$, $\mu_{1,k}=\rho_{1}e^{-z_{0,k}}$, $\mu_{2,k}=\rho_{2}e^{-z_{0,k}}$ where $\rho_{1}$ and $\rho_{2}$ is the minimum and maximum thrust magnitudes, respectively, $\alpha$ is the fuel depletion parameter, and $\Delta t$ is the time step between nodes. The matrices $A_{k}$ and $B_{k}$, and the vector $g$ in the dynamics are given by:
$\begin{split}A_{k}=\begin{bmatrix}I_{3}&I_{3}\Delta t\\ {\bf{0}}_{3\times 3}&I_{3}\end{bmatrix},\quad B_{k}=\begin{bmatrix}(\frac{1}{2}\Delta t^{2})I_{3}\\ I_{3}\Delta t\end{bmatrix}\\ g=\begin{bmatrix}0&0&-\frac{1}{2}g_{0}\Delta t^{2}&0&0&-g_{0}\Delta t\end{bmatrix}^{T}\end{split}$ (73)
where $I_{3}$ denotes $\times3$ identity matrix, $g_{0}$ represents the gravitational acceleration on the Martian surface, and the following parameter values were used in this numerical experiment.
@@
$\centering\begin{split}x_{\text{init}}=[200.0,\ 0.0,\ 800.0,\ -35.0,\ 0.0,\ -75.0]^{T}\\ \rho_{1}=7440,\ \rho_{2}=18600,\ g_{0}=3.7114,\ \theta_{\max}=\pi/12\\ m_{\text{dry}}=1505,\ m_{\text{wet}}=1905,\ \alpha=4.53\times 10^{-4}\end{split}\@add@centering$
Moreover, to validate the infeasibility detection capability of the developed solver, numerical experiments were conducted with both a feasible and an infeasible case. Since the Mars landing problem becomes infeasible if the flight time $t_{f}$ is shorter than the minimum admissible value [1, 22], we selected $t_{f}=48$s as the feasible case and $t_{f}=25$s as the infeasible case. The feasibility of these cases was confirmed by examining the solutions returned by ECOS. For both cases, ten problem instances were generated by increasing the number of nodes $N$ from 25 to 500. We set the termination tolerance to $10^{-8}$ for the developed customized solver and ECOS, while a relaxed tolerance of $10^{-4}$ was used for SCS, considering the convergence characteristics of ADMM.
TABLE I: Differences in Solutions for Feasible Mars Landing Problems
Differences Min. Max. Mean Median
$\|x^{\ast}-x^{\ast}_{\text{ECOS}}\|_{\infty}$ 7.899e-07 1.126e-01 5.325e-02 6.759e-02
$\|x^{\ast}-x^{\ast}_{\text{SCS}}\|_{\infty}$ 2.848e+00 1.344e+01 8.470e+00 8.783e+00
$\|x^{\ast}_{\text{ECOS}}-x^{\ast}_{\text{SCS}}\|_{\infty}$ 2.925e+00 1.343e+01 8.510e+00 8.819e+00
TABLE II: Residuals of the Developed Solver’s Solution
for Feasible Mars Landing Problems
Residuals Min. Max. Mean Median
$\|Ax^{\ast}-b\|_{\infty}$ 5.733e-12 3.621e-11 1.734e-11 1.425e-11
$\|Gx^{\ast}+s^{\ast}-h\|_{\infty}$ 4.489e-11 7.645e-05 1.457e-05 5.616e-08
$s^{\ast T}z^{\ast}$ 5.228e-10 2.862e-08 1.655e-08 1.825e-08
Results—Table I presents the differences in the converged solutions among the solvers for the feasible problem instances, where $x^{\ast}$, $x^{\ast}_{\text{ECOS}}$, and $x^{\ast}_{\text{SCS}}$ denote the solutions obtained by the developed solver, ECOS, and SCS, respectively. When compared with ECOS, the developed solver produced solutions that differed by values on the order of $10^{-7}$ to $10^{-1}$. These small discrepancies indicate that the developed solver produces accurate solutions consistent with those obtained by the existing solver. On the other hand, relatively larger discrepancies on the order of $10^{0}$-$10^{1}$ were observed for SCS, since it used a relaxed termination tolerance to accommodate its convergence characteristics.
The solution accuracy is further assessed through the residual values for feasibility and optimality as shown in Table II. Specifically, $\|Ax^{\ast}-b\|_{\infty}$ and $\|Gx^{\ast}+s^{\ast}-h\|_{\infty}$ quantify the maximum feasibility residuals for the equality and inequality constraints, respectively, while $s^{\ast T}z^{\ast}$ represents the duality gap, serving as a measure of optimality. Here, $A$, $b$, and $G$, $h$ denote the matrices and vectors corresponding to the equality and inequality constraints, respectively, which appear when the problem is expressed in the form of $\mathcal{P}$. The results show that the residual values on the order of the prescribed tolerance verify the accuracy of the solutions obtained by the developed solver.
@@
Figure 5 presents the computation time results for the feasible Mars landing trajectory optimization problems. It can be observed that the developed solver and ECOS achieve faster onboard computation times than SCS even under tighter termination tolerances, demonstrating the effectiveness of IPMs for G&C applications. For small-scale instances ($N\leq 150$), ECOS and the developed solver exhibit comparable performance. However, for larger-scale instances, the solver developed in this study demonstrates faster performance. This improvement is attributed to the elimination of unnecessary logic and the enhanced efficiency of data memory management achieved through solver customization.
Lastly, Table III presents the results demonstrating the infeasibility detection capability of the developed solver. It shows the converged values of $\kappa^{\ast}$ and $\tau^{\ast}$ together with the number of solver iterations for the infeasible instances. As described in Section III-B, the condition $\kappa^{\ast}>0$ with $\tau^{\ast}=0$ provides a certificate of infeasibility for the given problem. The resulting values of $\kappa^{\ast}$ and $\tau^{\ast}$, which satisfy the infeasibility conditions within the numerical tolerance, confirm that the solver successfully detects infeasibility. Moreover, it is observed that the solver achieves convergence to a certificate of infeasibility in only 7–8 iterations. This implies that, even for infeasible instances, the algorithm converges within a finite number of iterations and explicitly provides infeasibility information, which can be exploited to establish contingency strategies in G&C applications under abnormal conditions.
Figure 5: Computation time results for the feasible Mars landing trajectory optimization examples
TABLE III: Infeasibility Detection Results for Infeasible
Mars Landing Problems
NN $\kappa^{\ast}$ $\tau^{\ast}$ $\kappa^{\ast}/\tau^{\ast}$ Iteration
25 1.742e-01 1.131e-08 1.541e+07 8
50 1.760e-01 1.168e-08 1.506e+07 8
75 1.742e-01 1.131e-08 1.732e+05 8
100 1.765e-01 9.943e-07 1.775e+05 7
125 1.762e-01 9.752e-07 1.807e+05 7
150 1.749e-01 9.090e-07 1.924e+05 7
200 1.709e-01 8.097e-07 2.111e+05 7
300 1.665e-01 7.352e-07 2.264e+05 7
400 1.624e-01 6.645e-07 2.444e+05 7
500 1.690e-01 8.418e-09 2.007e+07 7
@@ VI-B Quadcopter Model Predictive Control
$\begin{split}\underset{x,u}{\text{minimize}}&\quad(x_{N}-x_{r})^{T}Q_{N}(x_{N}-x_{r})\ \\ &\quad+\sum_{k=0}^{N-1}\left[(x_{k}-x_{r})^{T}Q(x_{k}-x_{r})+u_{k}^{T}Ru_{k}\right]\\ \text{subject to}&\quad x_{k+1}=A_{k}x_{k}+B_{k}u_{k}\quad\forall k\in\{0,1,\cdots,N-1\}\\ &\quad x_{0}=x_{\text{init}}\\ &\quad x_{\min}\leq x_{k}\leq x_{\max}\quad\forall k\in\{0,1,\cdots,N\}\\ &\quad u_{\min}\leq u_{k}\leq u_{\max}\quad\forall k\in\{0,1,\cdots,N-1\}\end{split}$ (74)
@@
[…]
@@
Results—Table IV presents the solution differences among the solvers for the quadcopter MPC problems. As observed in the results, the solution obtained by the developed solver differed from those of existing solvers by the order of $10^{-3}$-$10^{-5}$, validating its accuracy. This accuracy is further corroborated by the residual values which remain at the prescribed tolerance level, as shown in Table V.
Figure 6 shows the computation time results for the quadcopter MPC problems. The results demonstrate that the solver developed in this study outperforms the other solvers. Such performance is mainly attributed to the efficient algorithmic formulation that directly handles quadratic cost functions and the fast convergence properties of IPMs. This capability is particularly important in MPC applications, where the optimization problem should be solved repeatedly in real-time based on feedback from the updated states of the quadcopter.
TABLE IV: Differences in Solutions for Quadcopter MPC Problems
Differences Min. Max. Mean Median
$\|x^{\ast}-x^{\ast}_{\text{ECOS}}\|_{\infty}$ 6.152e-05 4.093e-03 1.007e-03 6.320e-04
$\|x^{\ast}-x^{\ast}_{\text{SCS}}\|_{\infty}$ 1.983e-04 4.848e-03 1.295e-03 9.675e-04
$\|x^{\ast}_{\text{ECOS}}-x^{\ast}_{\text{SCS}}\|_{\infty}$ 1.091e-04 1.805e-03 1.054e-03 1.076e-03
TABLE V: Residuals of the Developed Solver’s Solution
for Quadcopter MPC Problems
Residuals Min. Max. Mean Median
$\|Ax^{\ast}-b\|_{\infty}$ 5.965e-09 1.079e-07 2.133e-08 6.698e-09
$\|Gx^{\ast}+s^{\ast}-h\|_{\infty}$ 8.882e-16 2.282e-09 9.309e-10 6.729e-10
$s^{\ast T}z^{\ast}$ 7.540e-06 9.970e-05 2.559e-05 1.304e-05
@@ VII Conclusion
In this work, we have developed a customized IPM-based solver specifically tailored for embedded real-time convex optimization. The developed solver is capable of handling problems within the class of SOCPs, while supporting infeasibility detection through a homogeneous embedding formulation. Unlike conventional homogeneous self-dual embedding approaches, which are limited to problems with linear cost functions, we adopt a homogeneous embedding framework for MCPs from previous work which can directly accommodate quadratic objective terms. This formulation enables efficient and robust algorithm design by avoiding the reformulation of quadratic objectives into conic constraints. To support embedded system deployment, the solver is implemented in C with complete static memory allocation, making it suitable for resource-constrained systems. A code generation tool is also developed to automate the solver customization process for a given problem family, exploiting structural sparsity offline and minimizing computation overhead during online execution. Numerical benchmark results demonstrate that the developed solver outperforms existing IPM- and ADMM-based solvers on small- to medium-scale problems in terms of computation time and reliability. Although the developed solver shows performance degradation for some large-scale problems due to the limited scalability inherent in IPM algorithm, it still achieves better performance than the other IPM-based solvers due to its efficient algorithmic design. Moreover, the results on embedded hardware platform experiments highlight its effectiveness for onboard optimization in G&C applications such as MPC and trajectory optimization.
Figure 6: Computation time results for the quadcopter MPC examples
@@ Acknowledgments
This work was partly supported by Theater Defense Research Center funded by Defense Acquisition Program Administration under Grant UD240002SD, and by the Korea Aerospace Research Institute (KARI)’s own research project titled “Development of Control Performance Analysis Program for Spaceplanes”. The authors acknowledge the use of ChatGPT to improve the syntax and grammar of several paragraphs in the manuscript. The manuscript was originally drafted by the authors, with the AI system employed to enhance clarity and readability without modifying the technical content.